Alex and Boy working together can complete a job in 7 1\2 hours. Working alone, Alex takes 8 hours longer than Boy to do the job. How long would it take each working alone?helllpp me...
Answer by edjones(7311)

b=hours it takes boy to do the job.
b+8=hours it takes Alex to do the job.
1/b=part of the job Boy can do in 1 hr.
1/(b+8)=part of the job Alex can do in 1 hr.
Together they can complete a job in 15/2 hours so they can do 2/15 of the job in one hour.
1/b + 1/(b+8) = 2/15
15(b+8)+15b=2b(b+8) multiply each side by 15b(b+8) to eliminate fractions.
15b+120+15b=2b^2+16b
2b^2+16b-15b-15b-120=0
2b^2-14b-120=0
2(b^2-7b-60)=0
(b-12)(b+5)=0
b=12 hrs it takes Boy working alone.
b+8=20 hrs it takes Alex working alone.
.
Ed